OpenStudy (anonymous):

Eliminate the parameter. x = t2 + 1, y = t2 - 1

OpenStudy (anonymous):

\[x-y\]

OpenStudy (anonymous):

x=y+2

OpenStudy (mertsj):

t is the parameter and from the first equation we see that t^2=x-1.

OpenStudy (mertsj):

So, replace t^2 in the second equation with x-1

OpenStudy (anonymous):

@Mertsj I get x=t^2+1, y=x-2 @tcarroll010 your're moving too fast

OpenStudy (mertsj):

Yes. Good job.

OpenStudy (anonymous):

@Mertsj and @tcarroll010 thank you both

OpenStudy (mertsj):

yw